find a polynomial function with the zeros -2 (multiplicity 2) and 2 (multiplicity 2), whose graph passes through the point (-3,125?
Answer by ewatrrr(24785)

f(x) = a(x+2)^2(x-2)^2
125 = a(-1)^2(-5)^2
5 = a
f(x) =5(x+2)^2(x-2)^2 = 5(x^4 -8x^2 + 16)